OpenStudy (anonymous):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the first equation is substituted into the second equation. x + 4y = –9 2x + 5y = –6 2x + 5(4y – 9) = –6 2x + 5(–4y – 9) = –6 2(4y – 9) + 5y = –6 2(–4y – 9) + 5y = –6

OpenStudy (anonymous):

explain?

OpenStudy (anonymous):

when you solve for x it would look like this x+4y=-9 -4y -4y x=-4y-9

OpenStudy (anonymous):

substitution just means making it all into one of the variables so when you solved for x you can plug that into your second equation. actually told you wrong sorry. 2(-4y-9)+5y=-6